I need to know how to solve this problem:
-3(3x-15)-(10+x)=35

Answer by checkley71(8403)

-3(3X-15)-(10+X)=35 OR -9X+45-10-X=35 OR -10X+35=35 OR -10X=0 OR X=0
PROOF -3(0-15)-(10-0)=35 OR 45-10=35 OR 35=35